# Information-theoretic machine learning for time-varying mode decomposition of separated airfoil wakes

### v1

@@ Abstract
We perform an information-theoretic mode decomposition for separated wakes around a wing. The current data-driven approach based on a neural network referred to as deep sigmoidal flow enables the extraction of an informative component from a given flow field snapshot with respect to a target variable at a future time stamp, thereby capturing the causality as a time-varying modal structure. We consider three examples of separated flows around a NACA0012 airfoil, namely, 1. laminar periodic wake at post-stall angles of attack, 2. strong vortex-airfoil interactions, and 3. a turbulent wake in a spanwise-periodic domain. The present approach reveals informative vortical structures associated with a time-varying lift response. For the periodic shedding cases, the informative structures vary in time corresponding to the fluctuation level from their mean values. With the second example of vortex-airfoil interactions, how the effect of vortex gust on a wing emerges in the lift response over time is identified in an interpretable manner. Furthermore, for the case of turbulent wake, the present model highlights structures near the wing and vortex cores as informative components based solely on the information metric without any prior knowledge of aerodynamics and length scales. This study provides causality-based insights into a range of unsteady aerodynamic problems.
@@ 1 Introduction
Along with the above modal-analysis techniques often used for cases with a time-invariant mean flow, there exists emerging data-driven and operator-based approaches that aim to be generalized for unsteady base flows. For example, a dynamic network-based data-driven approach known as broadcast mode analysis provides an input-output relationship through the construction of a complex network that describes the temporal evolution of vortical interactions in unsteady flows [13]. This technique identifies the structures that reflect the effect of perturbation in the network constructed by the flow snapshots. Optimally time-dependent mode decomposition has also been considered for a range of aerodynamic examples, which performs linear stability analysis for the instantaneously linearized dynamics [14]. Equipped with the time-dependent orthonormal modes derived through this linear technique, the dominant transient amplification of perturbations such as the perturbation dynamics about the time-varying mean flows can be analyzed [15, 16]. A variant of resolvent analysis applicable to time-evolving flows with sparsity promotion into the modal structures has been developed for the cases with time-varying mean velocity profiles [17]. By incorporating the temporal dimension into the numerical domain via a discrete time-differentiation operator, it is possible to examine linear amplification mechanisms for time-varying base flows as a change of the space–time modal structures. Furthermore, nonlinear machine-learning-based compression with autoencoder offers a low-rank manifold representation associated with nonlinear modes extracted from unsteady aerodynamic flow data [18, 19, 20].
@@
Aiming to capture the causal relationship between vortical structures and aerodynamic responses under either statistically stationary or time-varying mean flows, this study considers incorporating information theory into the data-driven identification of modal structures for unsteady aerodynamic flows. This is achieved by an information-theoretic data-driven approach, informative and non-informative decomposition [21]. For a given flow variable, this approach decomposes it into its informative and residual components with respect to a time-lagged target variable. Arranz and Lozano-Durán [21] reported that it is possible to extract the informative component from a given snapshot of turbulent channel flow through data-driven optimization, which maximizes the mutual information of the informative component with the wall-shear stress at a future time stamp while minimizing the mutual information with the residual component. With three examples of separated wakes around a NACA0012 airfoil, 1. laminar periodic flows at post-stall angles of attack, 2. strong vortex-airfoil interactions, and 3. a turbulent flow in a spanwise-periodic domain, the present study discusses how the information-theoretic approach, which has recently been used for some problems in turbulence [22, 23, 24, 25, 26, 27], characterizes the causality between the vortical flows and aerodynamic responses in a data-driven manner.
@@ 2 Methods
Figure 1: a) An example of the given flow field ${\bm{q}}$ and the informative component ${\bm{q}}_{I}$ extracted via the information-based mode extraction. b) Deep sigmoid flow used for the present decomposition.
@@
As illustrated in Fig. 1b and described in Eq. 9, the current model ${\cal F}$ takes the local coordinates $x$ and the lift response $C_{L}$ at a future time stump ${t+\Delta t}$ and provides the informative component $q_{I}$ at the current time $t$. To construct such an informative mode extractor ${\cal F}$, we use a fully-connected neural network (multi-layer perceptron) [33] with some constraints, refereed to as the deep sigmoidal flow [28]. This multi-layer perceptron is composed of bijective activation functions with positive weights. With these constraints, the deep sigmoidal flow achieves a bijective transformation, enabling the solution of ${q_{I}}$ produced by the mode extractor ${\cal F}$ to satisfy the constraint of the discounted Shannon information $H(C_{L}({\bm{x},t+\Delta t}),q_{I}({\bm{x}},t))=0$ in Eq. 5. The mutual information in this study is computed with the Gaussian kernel density estimator. The optimization in Eq. 10 is performed with the Adam algorithm [34]. Further details on the properties of deep sigmoidal flow with respect to the informative mode decomposition are referred to Arranz and Lozano-Durán [21].
@@ 3 Results
In this section, we apply a modified version of the approximated informative and non-informative mode decomposition, hereafter called informative mode decomposition or IMD for brevity, to three examples of separated airfoil wake, covering a range of spatiotemporal aerodynamic characteristics. The first example of laminar airfoil flow across the angles of attack exhibits periodic wake shedding while its frequency varies depending on the angle. The second case of vortex-gust airfoil interactions presents highly unsteady transient dynamics with a large excitation of aerodynamic forces during short time. Furthermore, the third problem of a turbulent separated wake involves the three-dimensionality of vortical structures with quasi-periodicity over time. Along with the current problem settings, let us examine how the present data-driven approach extracts physics for unsteady aerodynamic problems from an informatics point of view.
@@ 3.1 Example 1: periodic laminar airfoil wake shedding at post-stall angles of attack
Figure 2: Informative mode decomposition of periodic laminar airfoil wake shedding at post-stall angles of attack. Vorticity snapshots and resulting decomposed fields over phase $\theta$ across the angle of attack $\alpha$ are shown with the lift fluctuation $C_{L}^{\prime}$.
@@ 3.2 Example 2: strong vortex-gust airfoil interactions
Figure 3: Informative mode decomposition of strong vortex-airfoil interactions for positive gust with $G=1$ and $2$. Vorticity snapshots and resulting decomposed fields over time are shown with the lift coefficient $C_{L}$. The gray line in the lift response corresponds to the undisturbed case.
Since the case of $\alpha=40^{\circ}$ in the first example is taken as the baseline undisturbed condition, we use the same computational domain for the current simulation as that used for the first periodic shedding cases. The simulations with vortex gust encounters have also been verified and validated with previous studies [37, 18, 19]. Further details on the simulation setup are referred to as Fukami and Taira [18]. The baseline parameters of the convective time window $\Delta t$ and the weighting coefficient $\beta$ in the second example are set to 0.0085 and 0.01, respectively, although the dependence of the dynamic modal behavior on the parameter choice will be examined in Figs. 6 and 7.
@@
Figure 4: Informative mode decomposition of strong vortex-airfoil interactions for negative gust with $G=-1$ and $-2$. Vorticity snapshots and resulting decomposed fields over time are shown with the lift coefficient $C_{L}$. The gray line in the lift response corresponds to the undisturbed case.
@@
Figure 5: Dependence of informative modal structures on the weighting parameter $\beta$ of the cost function. Vorticity snapshots and resulting decomposed fields for $G=2$ over time are shown with the lift coefficient $C_{L}$. The gray line in the lift response corresponds to the undisturbed case.
@@
Figure 6: Dependence of informative modal structures on the convective time window $\Delta t$. Vorticity snapshots and resulting decomposed fields for $G=2$ over time are shown with the lift coefficient $C_{L}$. The gray line in the lift response corresponds to the undisturbed case.
@@ 3.3 Example 3: turbulent separated wake in a spanwise-periodic domain
Figure 7: Informative mode decomposition of turbulent separated wake at $Re=5,000$. Vorticity isocontours and resulting decomposed fields over time colored by spanwise velocity are shown with the lift coefficient $C_{L}$. The gray line in the lift response corresponds to the undisturbed case.
@@
Figure 8: Proper orthogonal decomposition for the LES data and a time series of the informative components (labelled as IMD) of the turbulent separated wake. a) The normalized amount of kinetic energy captured by POD modes. Inset: A snapshot of $\omega_{z}({\bm{x}},t)$. b) The first and second dominant POD modes extracted from the LES and informative fields.
@@ 4 Conclusions
This study performed an information-theoretic mode decomposition for separated wakes around a wing using a data-driven framework based on deep sigmoidal flow. The present approach extracts causally informative flow structures from an instantaneous snapshot with respect to their influence on future aerodynamic responses, enabling a data-driven time-varying modal analysis. Through three representative cases around a NACA0012 airfoil — periodic laminar wake, strong vortex-airfoil interactions, and turbulent wake — we showed that the method reveals informative vortical structures associated with the time-varying lift force. The current results highlight the potential of the present method to analyze causality-driven modal structures in complex unsteady aerodynamic flows.
The current mesh-free technique of deep sigmoidal flow was performed locally at each point in a given snapshot. It is not only adaptable but also computationally tractable for a range of aerodynamic scenarios. However, this local approach introduces some level of spatial discontinuities in the resulting modal structures, as also observed in the present study. Although performing mutual information-based optimization for the entire global field possesses challenges due to a large degree of freedom (the curse of dimensionality), the extension of the present formulation for convolutional neural network-based models may be promising for obtaining spatially continuous profiles in the mode while accounting for geometric flow features in the learning process. Furthermore, the theoretical connection between the present information-based technique and other data-driven approaches, such as Shapley additive explanations (SHAP) [52, 53], would be of interest to further deepen the understanding of data-driven causality analysis for unsteady flows. Equipped with these enhancements, the current nonlinear, information-theoretic framework may serve as a foundation for extracting causality-based insights into a range of unsteady aerodynamic problems.

### v2

# Information-theoretic machine learning for time-varying mode decomposition of separated aerodynamic flows

@@ Abstract
We perform an information-theoretic mode decomposition for separated aerodynamic flows. The current data-driven approach based on a neural network referred to as deep sigmoidal flow enables the extraction of an informative component from a given flow field snapshot with respect to a target variable at a future time stamp, thereby capturing the causality as a time-varying modal structure. We consider four examples of separated flows around a wing, namely, 1. laminar periodic wake at post-stall angles of attack, strong gust-wing interactions of 2. numerical and 3. experimental measurements, and 4. a turbulent wake in a spanwise-periodic domain. The present approach reveals informative vortical structures associated with a time-varying lift response. For the periodic shedding cases, the informative structures vary in time corresponding to the fluctuation level from their mean values. With the examples of gust-wing interactions, how the effect of gust on a wing emerges in the lift response over time is identified in an interpretable manner. Furthermore, for the case of turbulent wake, the present model highlights structures near the wing and vortex cores as informative components based solely on the information metric without any prior knowledge of aerodynamics and length scales. This study provides causality-based insights into a range of unsteady aerodynamic problems.
Nomenclature
$C_{L}$  = lift coefficient
$C_{L}^{\prime}$  = lift fluctuation
$c$  = chord length
$\cal F$  = informative mode extractor
$F_{L}$  = lift force
$G$  = gust ratio
$H$  = Shannon entropy
$I$  = mutual information
$p_{\bm{q}}$  = probability of the source variable
$p_{\bm{\lambda}}$  = probability of a target variable
$p_{{\bm{\lambda}},{\bm{q}}}$  = joint probability distribution of a target and the source variables
${q}$  = local flow field point
${q}_{I}$  = local informative component
${q}_{R}$  = local residual component
${\bm{q}}$  = global flow field
${\bm{q}}_{I}$  = global informative component
${\bm{q}}_{R}$  = global residual component
$R$  = state of the source variable
$R_{g}$  = radius of vortex gust
$\cal R$  = all the possible states of the source variable
$r$  = distance between the gust center and a point
$Re$  = Reynolds number
$S$  = state of a target variable
$\cal S$  = all the possible states of a target variable
$t$  = convective time
$u_{i}$  = velocity in the $i$th component
$u_{\infty}$  = free stream velocity
$u_{\theta}$  = rotational velocity
$w$  = spanwise velocity
${\bm{w}}$  = weights of the informative mode extractor
$\alpha$  = angle of attack
$\beta$  = weighting coefficient in cost function
$\omega$  = spanwise vorticity
$\nu$  = kinematic viscosity
$\rho$  = density
$\tau$  = wall-shear stress
$\theta$  = phase
${\bm{\lambda}}$  = target variable
${\Gamma}$  = circulation
${\Delta t}$  = time interval for causality estimation
${\Delta{\bm{x}}}$  = space interval for causality estimation
@@ 1 Introduction
Along with the above modal-analysis techniques often used for cases with a time-invariant mean flow, there exists emerging data-driven and operator-based approaches that aim to be generalized for unsteady base flows [13]. For example, a dynamic network-based data-driven approach known as broadcast mode analysis provides an input-output relationship through the construction of a complex network that describes the temporal evolution of vortical interactions in unsteady flows [14]. This technique identifies the structures that reflect the effect of perturbation in the network constructed by the flow snapshots. Optimally time-dependent mode decomposition has also been considered for a range of aerodynamic examples, which performs linear stability analysis for the instantaneously linearized dynamics [15]. Equipped with the time-dependent orthonormal modes derived through this linear technique, the dominant transient amplification of perturbations such as the perturbation dynamics about the time-varying mean flows can be analyzed [16, 17]. A variant of resolvent analysis applicable to time-evolving flows with sparsity promotion into the modal structures has been developed for the cases with time-varying mean velocity profiles [18]. By incorporating the temporal dimension into the numerical domain via a discrete time-differentiation operator, it is possible to examine linear amplification mechanisms for time-varying base flows as a change of the space–time modal structures. Furthermore, nonlinear machine-learning-based compression with autoencoder offers a low-rank manifold representation associated with nonlinear modes extracted from unsteady aerodynamic flow data [19, 20, 21].
@@
Aiming to capture the causal relationship between vortical structures and aerodynamic responses under either statistically stationary or time-varying mean flows, this study considers incorporating information theory into the data-driven identification of modal structures for unsteady aerodynamic flows. This is achieved by an information-theoretic data-driven approach, informative and non-informative decomposition [22]. For a given flow variable, this approach decomposes it into its informative and residual components with respect to a time-lagged target variable. Arranz and Lozano-Durán [22] reported that it is possible to extract the informative component from a given snapshot of turbulent channel flow through data-driven optimization, which maximizes the mutual information of the informative component with the wall-shear stress at a future time stamp while minimizing the mutual information with the residual component. With four examples of separated wakes around a wing, 1. laminar periodic flows at post-stall angles of attack, 2. numerically- and 3. experimentally-measured strong gust-wing interactions, and 4. a turbulent flow in a spanwise-periodic domain, the present study discusses how the information-theoretic approach, which has recently been used for some problems in turbulence [23, 24, 25, 26, 27, 28], characterizes the causality between the vortical flows and aerodynamic responses in a data-driven manner.
@@ 2 Methods
Figure 1: a) An example of the given flow field ${\bm{q}}$ and the informative component ${\bm{q}}_{I}$. b) Deep sigmoid flow.
@@
As illustrated in Fig. 1b and described in Eq. 9, the current model ${\cal F}$ takes the local coordinates $x$ and the lift response $C_{L}$ at a future time stamp ${t+\Delta t}$ and provides the informative component $q_{I}$ at the current time $t$. To construct such an informative mode extractor ${\cal F}$, we use a fully-connected neural network (multi-layer perceptron) [34] with some constraints, refereed to as the deep sigmoidal flow [29]. This multi-layer perceptron is composed of bijective activation functions with positive weights. With these constraints, the deep sigmoidal flow achieves a bijective transformation, enabling the solution of ${q_{I}}$ produced by the mode extractor ${\cal F}$ to satisfy the constraint of the discounted Shannon information $H(C_{L}({\bm{x},t+\Delta t}),q_{I}({\bm{x}},t))=0$ in Eq. 5. The mutual information in this study is computed with the Gaussian kernel density estimator. The optimization in Eq. 10 is performed with the Adam algorithm [35]. Further details on the properties of deep sigmoidal flow with respect to the informative mode decomposition are referred to Arranz and Lozano-Durán [22].
@@ 3 Results
In this section, we apply a modified version of the approximated informative and non-informative mode decomposition, hereafter called informative mode decomposition or IMD for brevity, to four examples of separated flows, covering a range of spatiotemporal aerodynamic characteristics. The first example of laminar airfoil flow across the angles of attack exhibits periodic wake shedding while its frequency varies depending on the angle. The second case of numerically-simulated vortex-gust airfoil interactions presents highly unsteady transient dynamics with a large excitation of aerodynamic forces during short time. While the third case of experimental data for strongly-disturbed flow around a wing also exhibits transient dynamics, these data include measurement noise. Furthermore, the fourth problem of a turbulent separated wake involves the three-dimensionality of vortical structures with quasi-periodicity over time. Along with the current problem settings, let us examine how the present data-driven approach extracts physics for unsteady aerodynamic problems from an informatics point of view.
@@ 3.1 Example 1: periodic laminar airfoil wake shedding at post-stall angles of attack
Figure 2: Informative mode decomposition of laminar airfoil wake. Vorticity snapshots and decomposed fields over phase across the angle of attack are shown with the lift fluctuation.
@@ 3.2 Example 2: strong vortex-gust airfoil interactions
Since the case of $\alpha=40^{\circ}$ in the first example is taken as the baseline undisturbed condition, we use the same computational domain for the current simulation as that used for the first periodic shedding cases. The simulations with vortex gust encounters have also been verified and validated with previous studies [38, 19, 20]. Further details on the simulation setup are referred to as Fukami and Taira [19]. The baseline parameters of the convective time window $\Delta t$ and the weighting coefficient $\beta$ in the second example are set to 0.0085 and 0.01, respectively, although the dependence of the dynamic modal behavior on the parameter choice will be examined later.
@@
Figure 3: Informative mode decomposition of strong vortex-airfoil interactions for positive gust. Vorticity and decomposed fields are shown with the lift coefficient (gray line: the undisturbed case).
@@
Figure 4: Informative mode decomposition of strong vortex-airfoil interactions for negative gust. Vorticity and decomposed fields are shown with the lift coefficient (gray line: the undisturbed case).
@@
Figure 5: Dependence of informative modal structures on the weighting parameter. Vorticity and decomposed fields are shown with the lift coefficient (gray line: the undisturbed case).
@@
Figure 6: Dependence of informative modal structures on the time window. Vorticity and decomposed fields are shown with the lift coefficient (gray line: the undisturbed case).
@@
3.3 Example 3: experimental measurements of large-amplitude transverse gust encounter
Figure 7: Informative mode decomposition of experimentally-measured large-amplitude transverse gust encounter at $Re=20,000$. Vorticity fields and resulting decomposed fields are shown with the lift coefficient.
To examine the robustness of the present approach against experimental flow measurements at a high Reynolds number, this section considers a flow around a flat-plate wing at a constant towing speed and angle of attack $\alpha$ passing through a large-amplitude transverse gust at a gust ratio $G=1.5$. The data set is made available by Towne et al. [48]. Experiments were performed in the University of Maryland free-surface water towing tank, and a flat-plate wing made of glass was used in measuring flow fields by particle image velocimetry. The time trace of aerodynamic forces is also collected along with that of flow fields, as shown in Fig. 7. The flat plate here is set to a fixed $\alpha=0^{\circ}$ and encounters a transverse gust with a sine-squared profile. The chord-length-based Reynolds number is 20,000. Both the vorticity field and the lift coefficient used in the present study are ensemble-averaged over eight and five individual runs, respectively. Further details on the experimental setups are referred to previous studies [49, 50, 48]. With the present strong gust ratio $G\geq 1$, this problem setup enables us to analyze the applicability of the approach for highly unsteady transient dynamic data measured under gusty and noisy conditions.
Let us perform the present informative mode decomposition for the time series of experimental measurements, as exhibited in Fig. 7. Here, the parameters of the convective time window $\Delta t$ and the weighting coefficient $\beta$ are set to 0.005 and 0.01, respectively. With the approach of the transverse gust, the lift coefficient sharply increases within sole two convective time and eventually is relaxed to the statistically steady state. Corresponding to this fast and large excitation of aerodynamic forces, the large-scale flow separation occurs due to the formation of a leading-edge vortex, as seen for $1\leq t\leq 1.74$. They eventually become turbulent fine structures and convect over a wing.
Similar to the numerical example of vortex-airfoil interactions, the gust does not appear in the informative modal field at the early stage of $t=0.250$ as the lift at that moment is primarily determined by the flow around a wing. Once the transverse gust begins to interact with a wing, the separated structures near the leading edge are extracted as the informative components. At $t=1.74$ presenting the lift peak, largely separated structures on the suction side and the newly-generated structure possessing negative vorticity near the trailing edge are highlighted in the informative modal field. Once the flow becomes turbulent and starts to dissipate at $t=3.06$, fine-scale structures with small contribution to the lift response are discarded from the informative mode. These observations indicate that structures deformed due to the gust-wing interaction mainly affect the lift dynamics, which coincides with our findings through the numerical example of strong vortex-airfoil interactions. This type of high Reynolds number aerodynamic flow can also be analyzed with the present informative mode decomposition.
Furthermore, noteworthy here is that the current approach offers time-varying informative modes while performing denoising. Such noisy components in experimental measurements often hinder the extraction of dominant structures in a data-driven manner, necessitating special care in the formulation such as sparsity promoting [51]. The successful extraction of clean modal structures indicates that the present information-theoretic framework finds the causal relationship between vortical structures and aerodynamic responses while removing noisy components that have no causal effects on lift dynamics, exhibiting the applicability to a range of aerodynamic flow scenarios.
@@ 3.4 Example 4: turbulent separated wake in a spanwise-periodic domain
Figure 8: Informative mode decomposition of turbulent separated wake at $Re=5,000$. Vorticity isocontours and decomposed fields colored by spanwise velocity are shown with the lift coefficient.
@@
Figure 9: POD for LES data and informative modes of the turbulent wake. a) Normalized kinetic energy. b) The first and second dominant POD modes.
@@ 4 Conclusions
This study performed an information-theoretic mode decomposition for separated wakes around a wing using a data-driven framework based on deep sigmoidal flow. The present approach extracts causally informative flow structures from an instantaneous snapshot with respect to their influence on future aerodynamic responses, enabling a data-driven time-varying modal analysis. Through four representative cases around a wing — periodic laminar wake, numerically- and experimentally-measured strong gust-wing interactions, and turbulent wake — we showed that the method reveals informative vortical structures associated with the time-varying lift force. The current results highlight the potential of the present method to analyze causality-driven modal structures in complex unsteady aerodynamic flows.
The current mesh-free technique of deep sigmoidal flow was performed locally at each point in a given snapshot. It is not only adaptable but also computationally tractable for a range of aerodynamic scenarios. However, this local approach introduces some level of spatial discontinuities in the resulting modal structures, as also observed in the present study. Although performing mutual information-based optimization for the entire global field possesses challenges due to a large degree of freedom (the curse of dimensionality), the extension of the present formulation for convolutional neural network-based models may be promising for obtaining spatially continuous profiles in the mode while accounting for geometric flow features in the learning process. Furthermore, the theoretical connection between the present information-based technique and other data-driven approaches, such as Shapley additive explanations (SHAP) [57, 58], would be of interest to further deepen the understanding of data-driven causality analysis for unsteady flows. As recently performed with the SHAP technique, one can consider performing reinforcement learning-based flow control based on the informative mode as a reward [59]. Examining the learned representation in the network may also help further assess what is learned as the causality between vortical structures and aerodynamic responses in an interpretable manner. Equipped with these enhancements, the current nonlinear, information-theoretic framework may serve as a foundation for extracting causality-based insights into a range of unsteady aerodynamic problems.